# Human-in-the-loop Learning for Dynamic Congestion Games

### v1

@@ I Introduction
  • Human-in-the-loop learning for dynamic congestion games: Prior studies focus on one-shot congestion games without information learning, while we study how users learn and alter traffic conditions on stochastic paths in a human-in-the-loop manner. When such congestion games meet human-in-the-loop learning, more users’ routing on stochastic paths generates both positive learning benefits and negative congestion effects there for followers. We use users’ positive learning effect to negate negative congestion, which fundamentally extends traditional one-shot congestion games (e.g., [20, 22] and [33]).
@@ II-A Dynamic Congestion Model
where $\alpha_{i}(t)$ is the correlation coefficient, measuring the leftover flow to be served over time. Note that $f(\cdot)$ is defined as a general increasing function in $\ell_{i}(t),n_{i}(t)$ and $\alpha_{i}(t)$.
@@ II-B Human-in-the-loop Learning Model
However, the fused summary $y_{i}(t)$ by $n_{i}(t)$ users can still be inaccurate. Given $n_{i}(t)$, we define the group probabilities under $\alpha_{i}(t)_{H}$ and $\alpha_{i}(t)=\alpha_{L}$ for observing a hazard below:
@@
  • At the beginning of time $t$, the crowdsourcing platform publishes latest latency $\mathbb{E}[\ell_{i}(t)|x_{i}(t-1),y_{i}(t-1)]$ and prior hazard belief $x_{i}(t)$ in (4) on any stochastic path $i$.
@@ III Problem Formulations for Myopic and Socially Optimal Policies
Based on the dynamic congestion and human-in-the-loop learning models in Section II, we next formulate the optimization problems for both myopic and socially optimal policies. The myopic policy is widely used in existing crowdsourcing platforms (e.g., Google Maps and Waze), defined as follows.
@@ III-B Problem Formulation for Socially Optimal Policy
Different from the myopic policy, the socially optimal policy minimizes the long-term social travel cost for all users. Define ${n_{i}^{*}(t)\in\{0,1,\cdots,N(t)\}}$ to be the optimal exploration number, and let vector $\mathbf{n}^{*}(t)$ summarize all the $n_{i}^{*}(t)$. Different from the myopic policy that only minimizes each user’s travel cost, the social optimum aims to well control the overall congestion cost and information learning to minimize the long-run expected social cost.
@@ IV Policies Comparison via PoA Analysis
In this section, we first analyze the monotonicity of the two policies with respect to hazard belief ${x_{i}(t)}$ and variance cost ${V(\cdot)}$. Then, we show that the myopic policy misses both exploration and exploitation over time, as compared to the social optimum, leading to $\text{PoA}\geq 2$. This implies at least doubled total travel cost for all users and motivates our mechanism design in Section V.

### v2

@@ I Introduction
  • Human-in-the-loop learning for dynamic congestion games: Prior studies focus on one-shot congestion games without information learning, while we study how users learn and alter traffic conditions on stochastic paths in a human-in-the-loop manner. We consider a typical multi-path parallel network to regulate selfish user arrivals’ information learning for dynamic congestion games. When such congestion games meet human-in-the-loop learning, more users’ routing on stochastic paths generates both positive learning benefits and negative congestion effects there for followers. We utilize users’ positive learning effect to negate negative congestion, which fundamentally extends traditional one-shot congestion games (e.g., [20, 22] and [33]).
@@ II-A Dynamic Congestion Model
where $\alpha_{i}(t)$ is the correlation coefficient, measuring the leftover flow to be served over time. Note that correlation function $f(\cdot)$ is defined as a general increasing function in $\ell_{i}(t),n_{i}(t)$ and $\alpha_{i}(t)$.
@@ II-B Human-in-the-loop Learning Model
However, the fused summary $y_{i}(t)$ by $n_{i}(t)$ users can still be inaccurate. Given $n_{i}(t)$, we define the group probabilities under $\alpha_{i}(t)=\alpha_{H}$ and $\alpha_{i}(t)=\alpha_{L}$ for observing a hazard below:
@@
  • At the beginning of each time slot $t$, the crowdsourcing platform publishes latest latency $\mathbb{E}[\ell_{i}(t)|x_{i}(t-1),y_{i}(t-1)]$ and prior hazard belief $x_{i}(t)$ in (4) on any stochastic path $i$.
@@ III Problem Formulations for Myopic and Socially Optimal Policies
Based on the dynamic congestion and human-in-the-loop learning models in Section II, we next formulate the optimization problems for both myopic and socially optimal policies, respectively. Note that the myopic policy is widely used in existing crowdsourcing platforms (e.g., Google Maps and Waze), which is defined as follows.
@@
While the socially optimal policy minimizes the long-term social travel cost for all users.
@@ III-B Problem Formulation for Socially Optimal Policy
Define ${n_{i}^{*}(t)\in\{0,1,\cdots,N(t)\}}$ to be the exploration number under the socially optimal policy, and let vector $\mathbf{n}^{*}(t)$ summarize all the $n_{i}^{*}(t)$. Different from the myopic policy that only minimizes each user’s travel cost, the socially optimal policy aims to well control the overall congestion cost and information learning to minimize the long-run expected social cost function.
@@ IV Policies Comparison via PoA Analysis
In this section, we first analyze the monotonicity of the two policies with respect to hazard belief ${x_{i}(t)}$ and variance cost ${V(\cdot)}$. Then, we show that the myopic policy misses both exploration and exploitation over time, as compared to the socially optimal policy. Finally, we prove $\text{PoA}\geq 2$ caused by the myopic policy, which implies at least doubled total travel cost for all users and motivates our mechanism design in Section V.